a table top is shaped like a rectangle whose perimeter is 46 in. the length is 3 in more than the width
what is the length and the width?

Answer by Cromlix(4381)

Width = x
Length = x + 3
Perimeter = 2*width + 2*length
46 = 2(x) + 2(x + 3)
46 = 2x + 2x + 6
46 = 4x + 6
4x = 46 - 6
4x = 40
x = 10
Width = 10ins
Length = 13ins
Hope this helps.
:-)